An excellent solution for cleaning grease stains from cloth or leather consists of the following: Carbon tetreachloride 80% (by volume), ligroin 16%, and amyl alcohol 4%. How many pints of each should be taken to make up 75 pints of solution?
Answer by jim_thompson5910(35256)

Carbon tetreachloride:

80% of 75 pints = 0.80*75 = 60

-------------------------------------------------------

Ligroin:

16% of 75 pints = 0.16*75 = 12

-------------------------------------------------------

Amyl Alcohol:

4% of 75 = 0.04*75 = 3

-------------------------------------------------------

Summary:

60 pints of Carbon tetreachloride
12 pints of Ligroin
3 pints of Amyl Alcohol
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------